# When Audits and Recounts Distract from Election Integrity: The 2020 U.S. Presidential Election in Georgia

### v1

@@ 3 First Count, Audit, and Recount Differ Substantially
Table 2: Election day, advance, absentee, and provisional vote tallies for Fulton County, GA, precinct RW01 in the 2020 U.S. Presidential election
@@
The hand count could easily be more accurate than the machine count. Indeed, it is well known that hand counts of hand-marked paper ballots are often more accurate than machine counts, in part because human readers can interpret faint, improper, and ambiguous marks better than machines can, even when the machines are working properly. Experience in Georgia in 2020 shows that Dominion’s scanner settings (low resolution, black-and-white) can cause voters’ selections not to appear at all in the images, selections that human readers looking at the actual ballots can easily discern. [^18] See, e.g., Judge Amy Totenberg’s Opinion and Order of 11 October 2020, in Curling et al. v. Raffensperger, 1:17-CV-2989-AT, at 4, 30, 95, 101, 103, 114–135.
Evidence that hand counts are more accurate than machine counts comes from recounts and studies of the “residual vote,” [4, 3, 5, 2, 1, 9] the number of invalid votes and undervotes. Manual tallies generally find more valid votes. Hand-count error rates are known to depend on many factors, including ballot design, the method for hand counting (“sort-and-stack” versus “read-and-mark”), and the size of counting teams. They presumably also depend on whether there are additional quality control measures in place, such as checking sorted piles of ballots to ensure that each pile has votes for just one candidate.
The study cited by the Georgia Secretary of State to support his claim of accuracy [11] is a laboratory study with 108 subjects and 120 ballots, each containing 27 contests with two candidates. It used three kinds of “ballots”: printout from two kinds of DRE (direct-recording electronic) voting system and an optical scan ballot. The highest error rates were for thermal printout from DREs, which does not resemble Georgia’s BMD printout nor Georgia’s hand-marked paper ballots. The method with the highest error was the “sort-and-stack” tally method that Georgia used in its audit. The study did not observe hand tabulation in a real election, nor did it involve BMD summary printout.
Differences between the original count and the machine recount are large and unexplained; for instance, the difference in the counts of Biden’s Absentee votes is almost 3 percent. It is now impossible to know what went wrong, nor whether the differences are primarily attributable to malware, bugs, misconfiguration, or human error.
@@ 3.1 The two machine counts in Fulton County
I examined the internal consistency of the two machine counts (the original machine count and the machine recount) in Fulton County using data from the election management system (EMS) including cast vote records (CVRs), scanned images of ballots, and BMD printout, and other files made available to the plaintiffs in Curling et al. v. Raffensperger et al. To confirm that the EMS data were the correct data, tallies were calculated and compared to the official results for Fulton County; they matched: [^19] First machine count results: https://results.enr.clarityelections.com/GA/Fulton/105430/web.264614/#/summary (visited 11 July 2024) Second machine count results: https://results.enr.clarityelections.com/GA/107231/web.264614/#/detail/5000?county=Fulton (visited 11 July 2024)
@@
Table 3: Data used to verify that the EMS download matched the official results in Fulton County, GA.
@@
The number of cast vote records in the two machine counts should be equal. Differences might occur if (i) some ballots or BMD printout cards were misplaced or found between the two machine counts, so a different number pieces of paper was scanned in the two machine counts; (ii) malware, bugs, misconfiguration, or a bad actor added, deleted, or altered records in the election management system in one or both machine counts; (iii) Fulton County did not scan every validly cast ballot or BMD printout card exactly once in each machine count. Compelling evidence that (ii) or (iii) is true is presented below, but all three could be true simultaneously.
@@
Entire batches of images are missing from Fulton County’s production, for example, images from Scanner 801 batch 117 and Scanner 801 batch 118 are referred to in the cast vote records for the second machine count but the images were not among the electronic records. Without additional information it is impossible to determine whether the missing images are missing because of human error, programming errors (bugs), or malware in Fulton County’s election management system (EMS)—possibilities that are not mutually exclusive.
It is nonetheless possible to use the images to show that Fulton County’s election results included many votes more than once in the reported tabulations. The full extent of multiple-counting cannot be determined without additional information, but there is evidence that it added thousands of bogus votes to the reported machine-count results. That is, thousands of Fulton County voters’ votes were included in the reported totals more than once. It is not possible to determine conclusively whether any voter’s votes were omitted from the reported totals.
@@
It is not generally possible to tell whether two 200dpi black-and-white scans of BMD printout cards are scans of the same piece of paper simply by looking at those two scans, because BMD printout cards containing the same votes look the same at low resolution in black-and-white. [^20] A sufficiently high-resolution scan might make it possible to identify differences in the arrangement of the paper fibers. See [10]. However, if both scans contain a rare write-in name or rare combination of write-in names, that is evidence of a duplicate. Similarly, if a series of votes is repeated in in the same order (or reverse order) in different scan batches of BMD printout, that is also evidence that they are repeated images of the same collection of paper. If the duplicated (or reversed) vote sequences are long and include rare write-in names, the evidence that they are scans of the same physical pieces of paper is compelling.
@@
Table 4: Images that were included in the original machine count in Fulton County at least twice. Images are posted at https://figshare.com/s/9819e969a8a6172c25bc (Appendix 2).
@@
Table 5: Images that were (erroneously) included in the machine recount at least three times. Images are posted at https://figshare.com/s/9819e969a8a6172c25bc (Appendix 3).
@@
Nine hundred sixteen (916) of the 2,871 sets of images were images of hand-marked paper ballots. In a random sample of 100 of those 916, I verified visually that 46 contained triplicate images. I confirmed the determination for 98 of the 100 sets. I disagreed about one set, and was unable to verify one set. Treating this conservatively as 98 agreements in 100 random checks yields a 95 percent lower confidence bound that at least 891 of the 918 claimed multiples are genuine multiples.
@@
These observations make it clear that Fulton County did not keep track of which ballots and BMD cards had been scanned and which had not, in both the original count and in the machine recount. It is also possible that the electronic records were altered accidentally or intentionally. The electronic records of the election are not intact. This is a surprising gap: the most basic election safeguard is to check whether the number of voters who participated is equal to the number of ballots and BMD printout cards that were cast and to the number that were tabulated. Moreover, One might reasonably expect all electronic election materials to be backed up onsite and offsite, at least for the federally mandated retention period of twenty-two months, so the loss of hundreds of thousands of image files from the first machine count and of nearly 18,000 images from the second machine count is hard to fathom.
@@
Fulton County would have noticed these errors if it kept track of ballots and BMD printout cards and checked the total number against the number reported in the electronic tabulation. It seems that Fulton County does not know how many ballots and BMD printout cards were cast in the election, how many voters cast votes, or how many pieces of paper were scanned—nor how those numbers compare to each other. Absent basic ballot accounting, pollbook reconciliation, and counting of electronic records, it is unsurprising that the two machine tallies differ so much. The U.S. Election Assistance Commission has published best practices for chain of custody. [^22] https://www.eac.gov/sites/default/files/bestpractices/Chain_of_Custody_Best_Practices.pdf accessed 11 July 2024
@@
Fulton County’s lax curation and processing of cast ballots, cast BMD printout, and electronic records make a true risk-limiting audit impossible because even a perfect tabulation of the votes from the available paper might not show who really won. Voters have good reason to believe that some votes counted more than others, since some votes were included twice or thrice in the totals. There is no way to know how many votes were omitted from the tabulation, absent access to the physical ballots and BMD printout and evidence that the chain of custody is intact. It is impossible to determine whether malware, bugs, misconfiguration, or malfeasance disenfranchised voters or altered the election results.
@@
The audit planning, process, and controls did not detect the double and triple counting. Even if Fulton County did not rely on ballot-marking devices for all in-person voters, the lack of basic accounting controls makes it impossible to determine who really won, even by hand counting the votes: the record of the vote could easily be incomplete or adulterated. There is no reason to believe that problems of the kinds described above are limited to Fulton County.
@@ 4 Summary
An accurate recount of the votes in a trustworthy paper trail can determine the true winners of an election, and a rigorous audit can provide confidence that a well-run election found the true winner(s). But neither a recount nor an audit can compensate for using untrustworthy technology to record votes or for a poorly run election with inadequate physical controls; in such circumstances, recounts and audits distract attention from the real problems rather than justifying confidence. Absent a trustworthy record of the votes, no procedure can provide affirmative evidence that the reported winner(s) really won.
Georgia lacks such a record for many reasons, including the heavy reliance on BMDs; lack of physical accounting of ballots, memory cards, and other election materials; lack of pollbook and voter participation reconciliation; lack of rigorous chain of custody; etc. To provide reasonable assurance that every validly cast vote is counted—accurately—requires systematic improvements:
@@
    Reduce the use of ballot-marking devices to a minimum:
    • BMDs do not necessarily print voters’ selections accurately. They can be hacked or misconfigured [12, 6].
    • A growing body of empirical work shows that few voters check the BMD printout, and those who do rarely catch and report errors [8, 14, 13].
    • There is no way for a voter to prove to an election official or anyone else that a BMD malfunctioned. Hence, there is no way to ensure that malfunctioning devices are removed from service if voters notice BMDs misbehaving. And if a device is caught misbehaving, there is no way to reconstruct the correct election outcome [6].
    • There is no way to test BMDs adequately prior to, during, or after an election to establish whether they altered votes, even if they altered enough votes to change electoral outcomes [6, 16].
@@
    Implement better procedures and checks on chain of custody of election materials, especially voted ballots. Georgia currently cannot determine whether every validly cast ballot was included in the reported results, whether there was electronic or physical “ballot-box stuffing,” or whether votes were altered. [^23] This is evidenced by the fact that the 2020 audit found thousands of untabulated ballots. Per the Secretary of State’s office, “[t]he audit process also led to counties catching making mistakes they made in their original count by not uploading all memory cards.” https://sos.ga.gov/news/historic-first-statewide-audit-paper-ballots-upholds-result-presidential-race accessed 11 July 2024. Because physical accounting for election materials was lacking, there is no way to know how many more votes validly cast in that election were not included in any of the reported tallies. Moreover, the lax recordkeeping evidently resulted in scanning the same batches of ballots more than once. Similarly, some ABBSs were presumably entered more than once, and as shown above, some were not entered at all.
@@
      Verify that the number of ballots sent to each polling location (and blank paper stock for ballot-marking devices and ballot-on-demand printers) equals the number returned voted, spoiled, or unvoted. This must be a physical check based on manual inventories, not on reports from the voting system.
@@
A genuine risk-limiting audit requires more than Georgia has attempted. Most importantly, it requires a demonstrably trustworthy record of voter intent. Georgia’s vote records are untrustworthy for many reasons, starting with the heavy use of ballot-marking devices, which do not produce a trustworthy record of the vote [6, 12], exacerbated by the lack of ballot accounting, pollbook reconciliation, and other elements of a good canvass. There are also problems with Georgia’s verification of voter eligibility and voter participation record. But even if every voter used a hand-marked paper ballot and there were no issues determining voter eligibility, Georgia simply does not keep track of election materials.

### v2

@@ 3 First Count, Audit, and Recount Differ Substantially
Tabelle 2: Election day, advance, absentee, and provisional vote tallies for Fulton County, GA, precinct RW01 in the 2020 U.S. Presidential election
@@
The hand count could easily be more accurate than the machine count. Indeed, it is well known that careful hand counts of hand-marked paper ballots are often more accurate than machine counts, in part because human readers can interpret faint, improper, and ambiguous marks better than machines can, even when the machines are working properly, as studies of “residual votes” and statewide recounts show [AR04, An18, AS05, ABS13, AAH13, Ca05]. [^16] Whether hand counts are more accurate than machine counts depends on many variables. The scrutiny and care involved in recounts and manual audits are generally higher than they are in initial hand counts. For instance, [An18] find that initial machine counts were often more accurate than initial hand counts—by using careful handcounts from statewide recounts as the touchstone for the correct counts.
The scanner settings Georgia uses for its Dominion scanners (low resolution, black-and-white) can cause voters’ selections not to appear at all in the images, selections that are obvious to human readers looking at the actual ballots. [^17] See, e.g., Judge Amy Totenberg’s Opinion and Order of 11 October 2020, in Curling et al. v. Raffensperger, 1:17-CV-2989-AT, at 4, 30, 95, 101, 103, 114–135. Manual tallies generally find more valid votes than machine tallies. Hand-count error rates are known to depend on many factors, including ballot design, the method for hand counting (“sort-and-stack” versus “read-and-mark”), and the size of counting teams. They presumably also depend on whether there are additional quality control measures in place, such as checking sorted piles of ballots to ensure that each pile has votes for just one candidate.
The study [GBG12] cited by Secretary Raffensperger to support his claim is a laboratory study with 108 subjects and 120 ballots, each containing 27 contests with two candidates. It used three kinds of “ballots”: printout from two kinds of DRE (direct-recording electronic) voting system and an optical scan ballot. The highest error rates were for thermal printout from DREs, which does not resemble Georgia’s BMD printout nor Georgia’s hand-marked paper ballots. The method with the highest error rate was the “sort-and-stack” tally method that Georgia used in its audit. The study did not observe hand tabulation in a real election, nor did it involve BMD summary printout.
The differences between the original count and the machine recount are large and unexplained; for instance, the difference in the counts of Biden’s Absentee votes is almost 3 percent. It is now impossible to know what went wrong, nor whether the differences are primarily attributable to malware, bugs, misconfiguration, or human error.
@@ 3.1 The two machine counts in Fulton County
This section assesses the internal consistency of the two machine counts (the original machine count and the machine recount) in Fulton County using data from the election management system (EMS) including cast vote records (CVRs), scanned images of ballots, and BMD printout, and other files made available to the plaintiffs in Curling et al. v. Raffensperger et al. To confirm that the EMS data were the correct data, tallies were calculated and compared to the official results for Fulton County; they matched: [^18] First machine count results: https://results.enr.clarityelections.com/GA/Fulton/105430/web.264614/#/summary (visited 11 July 2024) Second machine count results: https://results.enr.clarityelections.com/GA/107231/web.264614/#/detail/5000?county=Fulton (visited 11 July 2024)
@@
Tabelle 3: Data used to verify that the EMS download matched the official results in Fulton County, GA.
@@
The number of cast vote records in the two machine counts should be equal. Differences might occur if (i) some ballots or BMD printout cards were misplaced or found between the two machine counts, so a different number pieces of paper was scanned in the two machine counts; (ii) malware, bugs, misconfiguration, or a bad actor added, deleted, or altered records in the election management system in one or both machine counts; (iii) Fulton County did not scan every validly cast ballot or BMD printout card exactly once in each machine count; (iv) some scans were omitted or improperly included in one or both counts. Compelling evidence that (ii) or (iii) is true is presented below, but all four could be true simultaneously.
@@
Entire batches of images are missing from Fulton County’s production; for example, images from Scanner 801 batch 117 and Scanner 801 batch 118 are referred to in the cast vote records for the second machine count but the images were not among the electronic records. Without additional information it is impossible to determine whether the missing images are missing because of human error or malfeasance, programming errors (bugs), or malware in Fulton County’s election management system (EMS)—possibilities that are not mutually exclusive.
The extant images nonetheless prove that Fulton County’s election results included many votes more than once in the reported tabulations. The full extent of multiple-counting cannot be determined without additional information, but there is evidence that it added thousands of bogus votes to the reported machine-count results. That is, thousands of Fulton County voters’ votes were included in the reported totals more than once. It is not possible to determine conclusively whether any voter’s votes were omitted from the reported totals.
@@
It is not generally possible to tell whether two 200dpi black-and-white scans of BMD printout cards are scans of the same piece of paper simply by looking at those two scans, because BMD printout cards containing the same votes may be indistinguishable at low resolution in black-and-white. [^19] Differences in the monochrome threshold or scanner maintenance might create discernable differences. A sufficiently high-resolution scan might make it possible to identify differences in the arrangement of the paper fibers [Cl09]. However, if both scans contain a rare write-in name or rare combination of write-in names, that is evidence of a duplicate. Similarly, if a series of votes is repeated in in the same order (or reverse order) in different scan batches of BMD printout, that is also evidence that they are repeated images of the same collection of paper. If the duplicated (or reversed) vote sequences are long and include rare write-in names, the evidence that they are scans of the same physical pieces of paper is compelling.
@@
Tabelle 4: Images that were included in the original machine count in Fulton County at least twice. Images are posted at https://figshare.com/s/9819e969a8a6172c25bc (Appendix 2).
@@
Tabelle 5: Images that were (erroneously) included in the machine recount at least three times. Images are posted at https://figshare.com/s/9819e969a8a6172c25bc (Appendix 3).
@@
Nine hundred sixteen (916) of the 2,871 sets of images were images of hand-marked paper ballots. In a random sample of 100 of those 916, I verified visually that 46 contained triplicate images. I confirmed the determination for 98 of the 100 sets. I disagreed about one set, and was unable to verify one set. Treating this conservatively as 98 agreements in 100 random checks yields a 95 percent lower confidence bound that at least 891 of the 916 claimed multiples are genuine multiples.
@@
These observations make it clear that in the original count and in the machine recount, Fulton County did not keep track of which ballots and BMD cards had been scanned and which had not. It is also possible that the electronic records were altered accidentally or intentionally, or that some memory cards were not uploaded or uploaded more than once. The electronic records of the election are not intact. This is a surprising gap: the most basic election safeguard is to check whether the number of voters who participated is equal to the number of ballots and BMD printout cards that were cast and to the number that were tabulated. Moreover, one might reasonably expect all electronic election materials to be backed up onsite and offsite, at least for the U.S. federally mandated retention period of twenty-two months, so the loss of hundreds of thousands of image files from the first machine count and of nearly 18,000 images from the second machine count is hard to fathom.
@@
Fulton County would have noticed these errors if it had kept track of ballots and BMD printout cards and checked the total number against the number reported in the electronic tabulation. It seems that Fulton County did not know how many ballots and BMD printout cards were cast in the election, how many voters cast votes, or how many pieces of paper were scanned—nor how those numbers compare to each other. Absent basic ballot accounting, pollbook reconciliation, and counting of electronic records, it is unsurprising that the two machine tallies differ so much. The U.S. Election Assistance Commission has published best practices for chain of custody. [^21] https://www.eac.gov/sites/default/files/bestpractices/Chain_of_Custody_Best_Practices.pdf accessed 11 July 2024.
@@
Fulton County’s lax curation and processing of cast ballots, BMD printout, and electronic records make a true risk-limiting audit impossible because even a perfect tabulation of the votes from the available paper might not show who really won. Voters have good reason to believe that some votes counted more than others, since some votes were included twice or thrice in the totals. There is no way to know how many votes were omitted from the tabulation, absent access to the physical ballots and BMD printout and evidence that the chain of custody is intact. It is impossible to determine whether malware, bugs, misconfiguration, or malfeasance disenfranchised voters or altered the election results.
@@
The audit planning, process, and controls did not detect the double and triple counting. Even if Fulton County did not rely on ballot-marking devices for all in-person voters, the lack of basic accounting controls makes it impossible to determine who really won, even by a perfect hand count of the votes: the record of the vote could easily be incomplete or adulterated. There is no reason to believe that problems of the kinds described above are limited to Fulton County.
@@ 4 Summary
An accurate recount of the votes in a trustworthy record can determine the true winners of an election, and a rigorous audit can provide confidence that a well-run election found the true winner(s). But neither a recount nor an audit can compensate for using untrustworthy technology to record votes, for instance, because the election was run poorly and had inadequate physical security controls; in such circumstances, recounts and audits distract attention from the real problems rather than justifying confidence. Absent a trustworthy record of the votes, no procedure can provide affirmative evidence that the reported winner(s) really won.
Georgia lacks such a record for many reasons, including the heavy reliance on BMDs; lack of physical accounting of voted ballots, memory cards, and other election materials; lack of pollbook and voter participation reconciliation; lack of rigorous chain of custody; etc. To provide reasonable assurance that every validly cast vote is counted—accurately—requires systematic improvements:
@@
    Reduce the use of ballot-marking devices to a minimum: [^22] Hand-marked ballots should be offered to in-person voters by default, with access to a BMD available upon request. BMDs or other accessible means of marking a ballot should be set up in advance, so that it is available if and when a voter requests to use it. BMD printout should resemble hand-marked paper ballots to the extent possible, to preserve voter privacy: they should the same paper stock, have the same format as hand-marked paper ballots, and the marks should be printed to resemble hand-made marks, e.g., by digitizing actual hand-made marks.
    • BMDs do not necessarily print voters’ selections accurately. They can be hacked or misconfigured [Ha23, ADS20].
    • A growing body of empirical work shows that few voters check the BMD printout, and those who do rarely catch and report errors [Be20, KBW20, HI21].
    • There is no way for a voter to prove to an election official or anyone else that a BMD malfunctioned. Hence, there is no way to ensure that malfunctioning devices are removed from service if voters notice BMDs misbehaving. And if a device is caught misbehaving, there is no way to reconstruct the correct election outcome [ADS20].
    • There is no way to test BMDs adequately prior to, during, or after an election to establish whether they altered votes, even if they altered enough votes to change electoral outcomes [ADS20, SX22].
@@
    Implement better procedures and checks on chain of custody of election materials, especially voted ballots. Georgia currently cannot determine whether every validly cast ballot was included in the reported results exactly once, whether there was electronic or physical “ballot-box stuffing,” or whether votes were altered. [^23] This is evidenced by the fact that the 2020 audit found thousands of untabulated ballots. Per the Secretary of State’s office, “[t]he audit process also led to counties catching making mistakes they made in their original count by not uploading all memory cards.” https://sos.ga.gov/news/historic-first-statewide-audit-paper-ballots-upholds-result-presidential-race accessed 11 July 2024. Because physical accounting for election materials was lacking, there is no way to know how many more votes validly cast in that election were not included in any of the reported tallies. Moreover, the lax recordkeeping evidently resulted in scanning the same batches of ballots more than once. Similarly, some ABBSs were presumably entered more than once, and as shown above, some were not entered at all.
@@
      Verify that the number of ballots sent to each polling location (and blank paper stock for ballot-marking devices and ballot-on-demand printers) equals the number returned voted, spoiled, or unvoted. This check should be physical, based on manual inventories, not on reports from the voting system.
@@
A genuine risk-limiting audit requires a demonstrably trustworthy record of voter intent. Georgia’s vote records are untrustworthy for many reasons, starting with the heavy use of ballot-marking devices, which do not produce a trustworthy record of the vote [ADS20, Ha23] no matter how much logic and accuracy testing or election-day monitoring there is [SX22]. The lack of a trustworthy record is exacerbated in Georgia by the lack of ballot accounting, pollbook reconciliation, and other elements of a good canvass. There are also problems with Georgia’s verification of voter eligibility and voter participation record. But even if every voter used a hand-marked paper ballot and there were no issues determining voter eligibility, Georgia does not keep track of election materials adequately through physical inventories, custody logs, and other means.
@@
Literatur
  • [AAH13] Alvarez, R.M.; Atkeson, L.R.; Hall, T.E.: Evaluating Elections: A Handbook of Methods and Standards. Cambridge University Press, NY, 2013.
  • [ABS13] Alvarez, R.M.; Beckett, D.; Stewart, C.: Voting Technology, Vote-by-Mail, and Residual Votes in California, 1990–2010. Political Research Quarterly, 66(3):658–670, 2013.
  • [ADS20] Appel, A.W.; DeMillo, R.; Stark, P.B.: Ballot-marking devices cannot assure the will of the voters. Election Law Journal: Rules, Politics, and Policy, 19(3):432–450, 2020. Preprint: https://papers.ssrn.com/sol3/papers.cfm?abstract_id=3375755.
  • [An18] Ansolabehere, S.; Burden, B.C.; Mayer, K.R.; III, C. Stewart: Learning from Recounts. Election Law Journal, 17(2):100–116, 2018.
  • [AR04] Ansolabehere, S.; Reeves, A.: Using Recounts to Measure the Accuracy of Vote Tabulations: Evidence from New Hampshire Elections 1946–2002. In (Alvarez, R.M.; Hall, L.R. Atkesonand T.E., eds): Confirming Elections: Creating Confidence and Integrity Through Election Auditing. Palgrave MacMillan, NY, 2004.
  • [AS05] Ansolabehere, S.; Stewart, C.: Residual Votes Attributable to Technology. The Journal of Politics, 67(2):365–389, 2005.
  • [AS20] Appel, A.W.; Stark, P.B.: Evidence-Based Elections: Create a Meaningful Paper Trail, Then Audit. Georgetown Law Technology Review, 4.2:523–541, 2020. https://georgetownlawtechreview.org/wp-content/uploads/2020/07/4.2-p523-541-Appel-Stark.pdf.
  • [Be20] Bernhard, M.; McDonald, A.; Meng, H.; Hwa, J.; Bajaj, N.; Chang, K.; Halderman, J.A.: Can Voters Detect Malicious Manipulation of Ballot Marking Devices? In: 41st IEEE Symposium on Security and Privacy. IEEE, pp. 679–694, 2020.
  • [Ca05] Carrier, M.A.: Vote Counting, Technology, and Unintended Consequences. St. John’s Law Review, 79(3):645–687, 2005.
  • [Cl09] Clarkson, W.; Weyrich, T.; Finkelstein, A.; Heninger, N.; Halderman, J. A.; Felten, E. W.: Fingerprinting Blank Paper Using Commodity Scanners. 2009 30th IEEE Symposium on Security and Privacy, pp. 301–314, 2009.
  • [GBG12] Goggin, S.N.; Byrne, M.D.; Gilbert, J.E.: Post-Election Auditing: Effects of Procedure and Ballot Type on Manual Counting Accuracy, Efficiency, and Auditor Satisfaction and Confidence. Election Law Journal: Rules, Politics, and Policy, pp. 36–51, 2012.
  • [Ha23] Halderman, J.A.: Security Analysis of Georgia’s ImageCast X Ballot Marking Devices. https://storage.courtlistener.com/recap/gov.uscourts.gand.240678/gov.uscourts.gand.240678.1681.0.pdf, 2023. Expert report submitted in Curling v. Raffensperger, Civil Action No. 1:17-CV-2989-AT U.S. District Court for the Northern District of Georgia, Atlanta Division.
  • [HI21] Haynes, A.A.; III, M.V. Hood: Georgia Voter Verification Study. https://s3.documentcloud.org/documents/21017815/gvvs-report-11.pdf, 2021. last visited 31 October 2021.
  • [KBW20] Kortum, P.; Byrne, M.D.; Whitmore, J.: Voter Verification of BMD Ballots Is a Two-Part Question: Can They? Mostly, They Can. Do They? Mostly, They Don’t. Technical report, 2020.
  • [SW12] Stark, P.B.; Wagner, D.A.: Evidence-Based Elections. IEEE Security and Privacy, 10:33–41, 2012.
  • [SX22] Stark, P.B.; Xie, R.: They may look and look, yet not see: BMDs cannot be tested adequately. In (Krimmer, R.; Volkamer, M.; Duenas-Cid, D.; Kulyk, O.; Rønne, P.; Solvak, M.; Germann, M., eds): Proceedings of E-Vote-ID 2022, Lecture Notes in Computer Science, volume 13553, pp. 122–138. Springer-Nature, Cham, 2022.